### v2

# EvRT-DETR: Latent Space Adaptation of Image Detectors for Event-based Vision

Abstract
Event-based cameras (EBCs) have emerged as a bio-inspired alternative to traditional cameras, offering advantages in power efficiency, temporal resolution, and high dynamic range. However, the development of image analysis methods for EBCs is challenging due to the sparse and asynchronous nature of the data. This work addresses the problem of object detection for EBC cameras. The current approaches to EBC object detection focus on constructing complex data representations and rely on specialized architectures. We introduce I2EvDet (Image-to-Event Detection), a novel adaptation framework that bridges mainstream object detection with temporal event data processing. First, we demonstrate that a Real-Time DEtection TRansformer, or RT-DETR, a state-of-the-art natural image detector, trained on a simple image-like representation of the EBC data achieves performance comparable to specialized EBC methods. Next, as part of our framework, we develop an efficient adaptation technique that transforms image-based detectors into event-based detection models by modifying their frozen latent representation space through minimal architectural additions. The resulting EvRT-DETR model reaches state-of-the-art performance on the standard benchmark datasets Gen1 (mAP $+2.3$) and 1Mpx/Gen4 (mAP $+1.4$). These results demonstrate a fundamentally new approach to EBC object detection through principled adaptation of mainstream architectures, offering an efficient alternative with potential applications to other temporal visual domains. The code is available at: https://github.com/realtime-intelligence/evrt-detr.
1 Introduction
[…]
Figure 1: Object Detection Performance vs Inference Time. A summary of object detection performance (COCO mAP) vs inference time (ms) of various models on 1Mpx automotive dataset. Circle size is proportional to the number of model parameters. Inference times are reported for NVIDIA T4 GPU. Our EvRT-DETR models achieve state-of-the-art accuracy while maintaining competitive inference speeds.
Figure 2: Event-Based Camera Output Visualization. Sample frames from the 1Mpx dataset showing event polarity: red indicates positive events and blue indicates negative events.
These unique characteristics of EBC data – sparsity, asynchronicity, and temporal nature – create fundamental challenges when applying traditional computer vision techniques. Early attempts to directly apply conventional image-based object detectors to simple event representations yielded poor performance [26], highlighting the gap between these domains. This led to two predominant research directions: (1) constructing sophisticated event representations that transform EBC data into formats compatible with existing detectors (e.g., [39]), or (2) designing specialized architectures tailored to handle the temporal nature of event data (e.g., [24]). A common assumption in these approaches is that EBC data processing requires either specialized representations or architectures due to the fundamental differences between event-based and frame-based vision, leading to development paths that often diverge from mainstream computer vision research.
Meanwhile, mainstream object detection has seen remarkable progress, evolving from CNN-based architectures like YOLO [28] and Faster R-CNN [29] to transformer-based models like DETR [2]. Recent advances have culminated in highly efficient and accurate models such as the Real-Time DEtection TRansformer (RT-DETR) [36], which achieves exceptional performance while maintaining real-time inference speeds. Though primarily optimized for processing individual images, these state-of-the-art detectors offer powerful feature extraction capabilities that could potentially be leveraged for temporal data.
This evolution in mainstream detection creates an opportunity to bridge the gap between conventional and event-based vision. Rather than assuming fundamental incompatibility between these domains, we explore a complementary direction: adapting state-of-the-art frame-based detectors to effectively process event data with minimal modifications. In this work, we investigate how RT-DETR can be strategically adapted to the event-based domain, demonstrating that modern object detection architectures can successfully work on event data and opening a new research direction.
[…]
Building on this foundation, we introduce I2EvDet (Image-to-Event Detection), a technique that enables straightforward transformation of image-based object detectors into video-capable models. Our approach leverages latent space adaptation principles [16, 17] where we freeze the pre-trained RT-DETR model and insert a lightweight recurrent neural network (RNN) module within the encoder’s latent representation space. This design enables the model to capture temporal dynamics across frames while preserving the powerful representation capabilities of the original detector and adding minimal computational overhead. The resulting adaptation effectively leverages information across multiple time windows without requiring a specialized video detection architecture.
The combined approach, which we call EvRT-DETR, achieves state-of-the-art performance on standard EBC benchmarks, outperforming specialized methods on both Gen1 [6] (mAP $+2.3$) and 1Mpx [26] (mAP $+1.4$) datasets. Importantly, these results are achieved using only standard modules from natural image and video analysis, without requiring domain-specific modifications. This demonstrates not only the effectiveness of our approach for EBC data but also suggests a broader paradigm for adapting image detectors to temporal domains. As shown in Figure 1, our model achieves superior accuracy while maintaining competitive inference speeds, making it practical for real-world applications.
Contributions of this work:
  • We show that mainstream object detectors like RT-DETR, when trained on simple image-like representations of EBC data using standard procedures, can achieve performance comparable to specialized EBC methods, offering a complementary approach to domain-specific architectures.
  • We propose I2EvDet, an efficient adaptation technique that transforms image-based detectors into video-capable models through minimal architectural additions to the frozen pre-trained model, with potential applications to various temporal visual domains.
  • We validate our approach in the challenging domain of event-based cameras, where our EvRT-DETR achieves state-of-the-art performance on the standard Gen1 (mAP $+2.3$) and 1Mpx (mAP $+1.4$) EBC benchmarks while maintaining competitive inference speeds.
2 Related Work
Research at the intersection of event-based cameras and object detection spans multiple domains in computer vision. Current approaches primarily focus on two directions: designing efficient representations for the unique properties of event data, and developing specialized neural architectures to process these representations. However, a third direction - adapting mainstream computer vision models to event data - remains underexplored despite its success in other domains. This section reviews relevant work across these areas, highlighting how our approach bridges the gap between specialized event-based methods and mainstream object detection advances.
2.1 EBC Data Representations
An event camera produces a stream of event data in the form $(t,p,x,y)$, where $t$ is the timestamp of an event, $p$ is the polarity (positive or negative) of the brightness change, and $(x,y)$ is the location of the event pixel. To simplify the analysis, the event data often are transformed into alternative representations, more suitable for conventional image analysis algorithms.
The simplest image-like representations of event camera data are Event Frames or 2D Histograms [8], where events within fixed time windows are accumulated into 2D frames of shape $(2,H,W)$ with polarity as the first dimension and spatial coordinates as the remaining dimensions. A natural evolution is the Stacked 2D Histogram, which further partitions each frame into $T$ time intervals, creating a representation of shape $(2,T,H,W)$ that is typically reshaped to $(2T,H,W)$ and treated as a natural image. While these representations are straightforward, directly applying existing computer vision algorithms to them yielded poor performance in the past [26].
To achieve better performance, other representations are being explored. For example, a Time Surface (TS) representation [39, 8] is an image-like representation where pixel values encode the time elapsed since the last event in a given pixel. Thus, unlike the simplest fixed time window representations, TS potentially can encode arbitrarily distant past. Empirically, it has been shown that using the TS representations offers significantly better performance on the object detection tasks compared to the 2D Histogram representations [26].
Many other EBC data representations also have been proposed. The recent ERGO-12 [39] work has developed efficient image-like representations, achieving state-of-the-art object detection performance on the Gen1 dataset. Meanwhile, other approaches reformulate object detection as a three-dimensional (3D) problem by treating events as points in spatial-temporal space or discretizing them into voxel grids [8]. While computationally heavier, these representations can preserve exact event timing, potentially enabling better handling of overlapping objects and complex motions.
2.2 EBC Object Detection Architectures
Once the data representation is selected, one can start to develop object detection algorithms. There are several families of object detection methods designed for particular data representations.
The most straightforward approach applies conventional image detection methods directly to 2D Histogram frames, but this yields relatively poor performance [26]. More advanced methods like Recurrent Vision Transformer (RVT) [11] and State Space Vision Transformer (S5-ViT-B) [40] extend this approach by incorporating temporal memory to capture sequential information, significantly improving detection quality compared to frame-by-frame processing.
The most recent family of methods attempts to use more efficient event representations to further improve the object detection performance. For example, 12-channel Event Representation through Gromov-Wasserstein Optimization (ERGO-12) [39] constructs an optimized image-like event representation and shows that the standard object detection methods perform exceptionally well on it. Alternatively, Asynchronous Spatio-Temporal Memory Network (ASTMNet) [19] and Adaptive Event Conversion (AEC) [23] attempt to combine both better event representation and use novel neural architectures architecture to expand state-of-the-art results.
Multiple other methods have been developed either treating EBC object detection as a 3D detection problem [30, 26, 10], exploiting benefits of neuromorphic architectures [3], or exploring various hybrid approaches [24, 25].
Figure 3: Overview of the I2EvDet framework applied to RT-DETR. Our approach adapts a pre-trained RT-DETR model to process temporal event data through minimal architectural modifications. The backbone and encoder components remain frozen during temporal adaptation, while trainable temporal modules are strategically inserted at multiple feature scales. Left: Event camera frames are processed through the frozen RT-DETR backbone and encoder. Center: Our temporal adaptation modules operate on multiscale features through residual connections (red pathways), enabling information flow across time steps while preserving spatial representation integrity. Right: Feature fusion and decoder stages produce final detections. The detailed view of the Temporal Adaptation Module (bottom right) shows the projection operations ($W_{i}^{down},W_{i}^{up}$) and recurrent memory module ($\mathbf{R}_{i}$) that enable efficient temporal modeling with minimal additional parameters.
2.3 Transformer-based Object Detection
Object detection has been revolutionized by transformer architectures in recent years. DEtection TRansformer (DETR) [2] was the first widely successful Transformer-based [33] model for object detection, offering a simple and elegant alternative to traditional CNN-based architectures [12, 28].
DETR follows a hybrid CNN-Transformer design with a backbone-encoder-decoder structure. It uses a CNN backbone (typically ResNet [15]) for feature extraction, followed by a Transformer encoder that processes these features into tokens and captures their correlations. The Transformer decoder then identifies objects through cross-attention to the encoder outputs.
Despite its innovative approach, the original DETR suffered from limitations including training instability [34], poor performance on small objects [2], and high computational costs during inference [38]. Subsequent research has addressed these issues through architectural improvements such as Feature Pyramid Networks (FPN), specialized decoder tokens, and deformable attention mechanisms [38, 2].
These incremental advances culminated in the Real-Time DEtection TRansformer (RT-DETR) [36], which achieves exceptional performance across all object scales with stable training and significantly faster inference times. RT-DETR represents the current state-of-the-art in efficient transformer-based detection and serves as an ideal foundation for adaptation to specialized domains such as event-based vision.
2.4 Model Adaptation Techniques
While specialized architectures dominate event-based vision, adapting mainstream architectures offers compelling advantages. Parameter-efficient methods [16, 14, 17] modify pre-trained models with minimal changes by adding compact modules to frozen backbones. These approaches have proven effective across language [7], vision [37], and multimodal systems [13], but remain underexplored for event-based vision despite their efficiency and transfer learning benefits.
Given the success of adaptation techniques in other domains and the power of transformer-based detection models, there is a clear opportunity to bridge these approaches for event-based vision. Our work explores this intersection, demonstrating that mainstream object detectors can be effectively adapted to EBC data through minimal modifications rather than requiring specialized architectures.
3 Methods
This section presents I2EvDet (Image-to-Event Detection), our adaptation framework for event-based object detection. Unlike approaches that rely on specialized architectures or complex representations, I2EvDet bridges mainstream object detection with temporal event data through strategic, minimal modifications. Our framework demonstrates how image-based detectors can be efficiently transformed into video-capable models while preserving their core strengths. We first describe our image-like representation of event data and then introduce the I2EvDet framework. Finally, we detail the technical implementation and discuss how our approach addresses the unique temporal dependencies in event camera data.
3.1 Event Representation
Our adaptation framework begins with transforming EBC data into an image-like representation that’s compatible with mainstream object detection models. We employ the Stacked 2D Histogram representation [8] as it provides a straightforward bridge between event data and conventional image processing architectures.
[…]
To construct a frame corresponding to an interval $[t_{0},t_{0}+T_{\text{frame}}]$, we create an intermediate stacked histogram $S$ from a set of events $\mathcal{E}$ in that interval:
$S(p,t_{i},y,x)=\sum_{\left\{e|t_{e}\in[t_{0},t_{0}+T_{\text{frame}})\right\}}\delta_{x}^{x_{e}}\delta_{y}^{y_{e}}\delta_{p}^{p_{e}}\mathbf{1}_{[t_{i},t_{i+1})}(t_{e})$
where $t_{i}=t_{0}+i\cdot T_{\text{bin}}$, $i\in[0,10)$ is the bin index, $(x,y)$ are spatial indices of the event in the EBC matrix, and $\mathbf{1}$ is an indicator function. Once a stacked histogram $S$ of shape $(2,10,H,W)$ is constructed, we merge the polarity and time dimensions to obtain an image-like 2D frame $F$ of shape $(20,H,W)$.
3.2 I2EvDet: Image-to-Event Detection Framework
We introduce I2EvDet (Image-to-Event Detection), a principled framework for adapting image-based object detectors to effectively process temporal data. While specialized architectures are typically developed for event-based object detection, our approach demonstrates that strategic adaptation of mainstream detectors can achieve superior performance with minimal architectural modifications.
Our I2EvDet framework applies latent space adaptation principles [16, 14, 17], which enable domain adaptation through minimal changes to pre-trained models. Our framework follows a two-stage process:
First, we establish a strong foundation by training the RT-DETR model on individual EBC frames (subsection 3.1), treating them as conventional images. This creates a detector with powerful spatial representation capabilities optimized for event data.
In the second stage, we transform this image-based detector into a video-capable model. We freeze the pre-trained RT-DETR parameters and insert lightweight RNN modules into the latent space of the encoder. The resulting model, which we call EvRT-DETR (Event RT-DETR), captures temporal dynamics across frames while preserving the powerful representation capabilities of the original detector and adding minimal computational overhead.
This modular approach offers several methodological advantages. The architecture enforces an explicit disentanglement between spatial and temporal feature processing, facilitating a hierarchical optimization regime where each component operates within its disentangled parameter subspace. By constraining the RT-DETR parameters after initial convergence, we preserve the integrity of the learned latent manifold during temporal adaptation, mitigating potential representation collapse that could occur during end-to-end optimization. Moreover, this design maintains the architectural factorization inherent in RT-DETR, enabling computational flexibility through selective module deactivation when temporal reasoning is unnecessary for inference scenarios.
Figure 4: Temporal Adaptation Benefits of EvRT-DETR. Visualization of object detection across sequential 1Mpx frames captured by a vehicle-mounted camera that gradually comes to a stop. While the camera is in motion (left frames), both detectors maintain accurate bounding boxes of stationary vehicles. As camera motion decreases (middle frames), event data becomes increasingly sparse. When the camera stops completely (right frames), RT-DETR (top row) loses detection capability while EvRT-DETR (bottom row) maintains consistent detection by leveraging its temporal memory module. This demonstrates how our temporal adaptation effectively addresses a fundamental limitation of image object detection methods when applied to event-based cameras.
3.3 Temporal Modeling Considerations
A key theoretical consideration in our adaptation framework is the choice of temporal memory mechanism. Event cameras present a unique challenge: static objects become “invisible” as they generate no events when motionless. Their detection requires having a persistent memory of past events. While Transformers have dominated recent vision architectures, their fixed context windows limit temporal memory. We selected RNNs for their theoretically unbounded memory capacity through recurrent state, making them uniquely suited for maintaining object persistence during periods of invisibility in the event stream. This theoretical foundation informed our EvRT-DETR implementation..
3.4 Technical Implementation of EvRT-DETR
Our EvRT-DETR architecture implements the I2EvDet framework through strategic, minimal modifications to the RT-DETR model. As shown in Figure 3, we identify key integration points within the pre-trained model’s latent space where temporal information can be effectively incorporated without disrupting spatial feature extraction.
The RT-DETR encoder produces multiscale feature representations $\left\{\mathcal{E}_{3},\mathcal{E}_{4},\mathcal{E}_{5}\right\}$, where each element corresponds to a Transformer token encoding rich spatial information. These features, originating from the last three stages of the backbone $\left\{\mathcal{S}_{3},\mathcal{S}_{4},\mathcal{S}_{5}\right\}$ [36], provide an ideal latent manifold for temporal adaptation.
Following latent space adaptation principles, we introduce specialized temporal processing modules at each scale, implementing three parallel RNN-based memory modules $\{\mathbf{R}_{3},\mathbf{R}_{4},\mathbf{R}_{5}\}$. This multi-scale temporal adaptation design ensures that temporal dynamics are captured at different levels of feature abstraction. Critically, each RNN module interfaces with its corresponding encoder feature map through residual connections, mathematically expressed as:
$\mathcal{E}_{i}^{t,proj}$ $=W_{i}^{down}\cdot\mathcal{E}_{i}^{t}$ (1)
$\left(\mathcal{O}_{i}^{t,proj},\mathcal{M}_{i}^{t}\right)$ $=\mathbf{R}_{i}\left(\mathcal{E}_{i}^{t,proj},\mathcal{M}_{i}^{t-1}\right)$ (2)
$\mathcal{O}_{i}^{t}$ $=W_{i}^{up}\cdot\mathcal{O}_{i}^{t,proj}$ (3)
~$\tilde{\mathcal{E}}_{i}^{t}$ $=\mathcal{E}_{i}^{t}+\alpha_{i}\cdot\mathcal{O}_{i}^{t},$ (4)
where $W_{i}^{down}\in\mathbb{R}^{C_{i}^{RNN}\times C_{i}}$ and $W_{i}^{up}\in\mathbb{R}^{C_{i}\times C_{i}^{RNN}}$ are the projection matrices that transform features into and out of the temporal processing space, $\mathcal{M}_{i}^{t}\in\mathbb{R}^{C_{i}^{RNN}\times H_{i}\times W_{i}}$ represents the temporal memory state for the $i$-th feature scale at time $t$, $\mathcal{O}_{i}^{t}$ denotes the corresponding temporal adaptation output, and $\alpha_{i}$ is a learnable scaling factor that modulates the influence of temporal information on the spatial representation, inspired by the ReZero technique [1]. In our implementation, the input projection matrix $W_{i}^{down}$ is integrated within the RNN architecture, which already handles dimensionality reduction through its internal gating mechanisms, while the output projection $W_{i}^{up}$ is explicitly implemented to ensure proper integration back into the frozen RT-DETR architecture. The transformed feature representation ~$\tilde{\mathcal{E}}_{i}^{t}$ thus incorporates temporal context while preserving the integrity of the original spatial representation $\mathcal{E}_{i}^{t}$, with our ablation studies showing that $C_{i}^{RNN}$ can be reduced 4x relative to $C_{i}$ with a minimal performance impact.
For the specific implementation of the RNN modules, we employ the ConvLSTM architecture [31], which effectively balances representational capacity with computational efficiency. This choice was informed by ConvLSTM’s demonstrated effectiveness in handling spatio-temporal data in the RVT model [11], allowing our adaptation to leverage established temporal modeling techniques.
4 Evaluation of the EvRT-DETR Model
Model Gen1 1Mpx/Gen4 Params (M)
mAP (%) Runtime (ms) mAP (%) Runtime (ms)
AsyNet [22] 14.5 - - - 11.4
AEGNN [30] 16.3 - - - 20.0
Spiking DenseNet [3] 18.9 - - - 8.2
Events-RetinaNet [26] 34.0 - 18.0 - 32.8
RED [26] 40.0 16.7 43.0 39.3 24.1
AEC [23] 47.0 10.6 (3.9) 48.4 37.6 (13.8) 46.5
RVT-B [11] 47.2 10.2 47.4 11.9 18.5
GET-T [24] 47.9 16.8 48.4 18.2 21.9
S5-ViT-B [40] 47.4 8.2 47.2 9.6 17.5
ASTMNet [19] 46.7 35.6 48.3 72.3 $>100$
SAST-CB [25] 48.2 - 48.7 57.5 (19.7) 18.9
ERGO-12 [39] 50.4 69.9 40.6 100.0 59.6
RT-DETR-T (ours) 46.0 6.4 44.1 8.6 20.1
RT-DETR-B (ours) 47.6 10.5 45.2 14.9 42.8
EvRT-DETR-T (ours) 52.3 8.4 49.9 12.5 34.4
EvRT-DETR-B (ours) $\mathbf{52.7}$ 12.7 $\mathbf{50.1}$ 18.8 57.1
Table 1: Performance Comparison on Event-Based Camera Datasets. Our approach achieves state-of-the-art results on both Gen1 and 1Mpx benchmarks, with significant improvements over previous methods. The first stage of our framework (RT-DETR-T/B) already achieves competitive performance using only image-based detection, while the proposed I2EvDet adapted models (EvRT-DETR-T/B) surpass all existing approaches. Runtime measurements are reported in milliseconds on an NVIDIA T4 GPU; values in parentheses represent original reported times on different hardware, with T4-equivalent times calculated using FLOPS/throughput conversion.
We evaluate our EvRT-DETR model on two standard EBC object detection benchmarks: Gen1 [6] and 1Mpx [26] (also known as Gen4) datasets. These datasets contain diverse driving scenarios captured by Prophesee’s neuromorphic vision sensors, providing an ideal testbed for validating our adaptation approach under real-world conditions.
4.1 Experimental Setting
Datasets.
We evaluate on two standard Prophesee automotive datasets: Gen1 [6] and 1Mpx [26]. Gen1 contains 39+ hours of recordings at $304\times 240$ resolution with manual annotations (1-4 Hz) for cars and pedestrians, presenting challenges with sparse labels and stationary objects. 1Mpx offers 14+ hours at higher resolution ($1280\times 720$) with denser automatic annotations (60 Hz) for cars, pedestrians, and two-wheelers. Both datasets are organized as 60-second clips with standard train/test splits. 1Mpx’s higher resolution and frame rate enables evaluation on more complex scenes, while Gen1’s annotation pattern tests robustness to sparse temporal supervision.
Event Representations.
[…]
For the Gen1 dataset (original size $(240,304)$), we pad frames to $(256,320)$ to ensure divisibility by 32, as required by the feature pyramid architecture of RT-DETR. For 1Mpx, we downsize from $(720,1280)$ to $(360,640)$ using bilinear interpolation and pad to $(384,640)$. This maintains consistency with prior work while allowing fair computational comparison.
Two-Stage Training Process.
We train EvRT-DETR in two stages: (1) training the RT-DETR detector on individual event frames using standard practices (Adam optimizer, EMA of weights), then (2) freezing these parameters and training only the lightweight RNN modules on both random and sequential clips to develop temporal memory. The first stage establishes strong spatial feature extraction, while the second adds temporal processing with minimal parameter overhead. For temporal training, we maintain RNN memory continuity across sequential clips while resetting for random clips, using different clip lengths for Gen1 (21 frames) and 1Mpx (10 frames). Full implementation details are provided in the supplementary material.
Evaluation Protocol
We evaluate EvRT-DETR using the standard COCO (Common Objects in Context) Mean Average Precision (mAP) measure [20], which allows direct comparison with existing methods. For consistency with previous works, we use the EBC-specific implementation of the COCO metrics provided by Prophesee’s Automotive Dataset Toolbox [27, 26]. All inference times reported in Table 2 are standardized to NVIDIA T4 GPU performance for fair comparison.
4.2 Performance Analysis
Table 1 presents a comprehensive comparison of our EvRT-DETR models against state-of-the-art approaches on the Gen1 and 1Mpx benchmarks. The results can be analyzed in terms of different methodological approaches to event-based object detection:
First, our baseline adaptation using two RT-DETR variants—RT-DETR-T (Tiny, with ResNet-18 backbone) and RT-DETR-B (Base, with ResNet-50 backbone) - trained on simple event frames achieves competitive performance ($46.0\%$ and $47.6\%$ mAP on Gen1, $44.1\%$ and $45.2\%$ mAP on 1Mpx) without requiring specialized architectures or complex event representations. This demonstrates that modern mainstream object detectors can be effectively adapted to event data through appropriate training.
More importantly, our complete EvRT-DETR models with temporal adaptation significantly outperform all existing methods, achieving new state-of-the-art results on both benchmarks. EvRT-DETR-B reaches $52.7\%$ mAP on Gen1 ($+2.3\%$ over the previous best) and $50.1\%$ mAP on 1Mpx ($+1.4\%$ over the previous best). Even our lighter EvRT-DETR-T model surpasses all existing approaches while maintaining faster inference times.
As shown in Figure 4, the temporal adaptation module provides crucial capabilities for handling stationary objects - a fundamental challenge for event cameras. When objects are in motion, both RT-DETR and EvRT-DETR perform well, but when motion stops and event generation becomes minimal, EvRT-DETRs temporal memory enables consistent detection.
The performance improvements come with minimal computational overhead. While we focus primarily on detection accuracy in this work, the runtime measurements in Table 1 demonstrate that our approach maintains competitive inference speeds despite adding temporal processing capabilities.
4.3 Analysis of Adaptation Design Choices
We systematically analyze key design choices in our temporal adaptation module to better understand the factors contributing to EvRT-DETR’s performance. This analysis provides insights into efficient adaptation strategies for event-based object detection.
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
EvRT-DETR-B (KS=1) $52.3$ $81.4$ $55.4$
EvRT-DETR-B (KS=3) $\mathbf{52.7}$ $\mathbf{82.0}$ $\mathbf{56.0}$
EvRT-DETR-B (KS=5) $52.0$ $81.3$ $54.8$
Table 2: Effect of ConvLSTM Kernel Size. Ablation study on Gen1 dataset showing optimal temporal adaptation performance with $3\times 3$ kernels
4.3.1 Spatial Context Size in Temporal Processing
While RVT [11] found optimal performance with $1\times 1$ ConvLSTM kernels (effectively pointwise LSTMs), our adaptation approach shows different optimal characteristics. As shown in Table 2, EvRT-DETR achieves best performance with $3\times 3$ kernels, suggesting that spatial context is valuable when adapting frozen RT-DETR features. This demonstrates how adaptation design choices may differ from specialized architectures built from the ground up. Increasing to $5\times 5$ kernels degrades performance, suggesting that the additional parameters and receptive field expansion do not provide beneficial information for the task given the available training data.
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
RT-DETR-B (ours) 47.6 76.3 49.5
EvRT-DETR-B $(0,1,1)$ $51.0$ $80.7$ $54.0$
EvRT-DETR-B $(1,0,1)$ $52.2$ $81.3$ $55.1$
EvRT-DETR-B $(1,1,0)$ $52.4$ $81.7$ $55.3$
EvRT-DETR-B $(1,1,1)$ $\mathbf{52.7}$ $\mathbf{82.0}$ $\mathbf{56.0}$
Table 3: Effect of Temporal Module Placement. Performance on Gen1 dataset with ConvLSTM modules at different feature scales. The array notation $(x,y,z)$ indicates presence (1) or absence (0) of temporal modules at scales $\mathcal{E}_{3}$, $\mathcal{E}_{4}$, and $\mathcal{E}_{5}$ respectively. Low-level features benefit most from temporal adaptation.
4.3.2 Multi-scale Temporal Adaptation
A key aspect of our approach is the insertion of temporal modules at multiple scales in the feature hierarchy. Table 3 shows performance when selectively removing ConvLSTM cells from specific feature levels, with the array notation $(x,y,z)$ indicating presence (1) or absence (0) at each scale $\{\mathcal{E}_{3},\mathcal{E}_{4},\mathcal{E}_{5}\}$.
The results reveal a clear pattern: the lowest-level features ($\mathcal{E}_{3}$) benefit most significantly from temporal adaptation, with performance dropping substantially when this module is removed. The impact diminishes at higher feature levels, suggesting that temporal information is most valuable for fine-grained, lower-level features that capture detailed motion patterns. This insight can guide more efficient adaptation designs where computational resources are directed to where they provide maximum benefit.
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%) $N_{\text{T}}$ (M)
EvRT-DETR-B (M=64) $52.1$ $81.3$ $54.9$ $2.3$
EvRT-DETR-B (M=128) $52.5$ $81.7$ $55.6$ $5.4$
EvRT-DETR-B (M=256) $52.7$ $\mathbf{82.0}$ $\mathbf{56.0}$ $14.4$
EvRT-DETR-B (M=512) $\mathbf{52.9}$ $81.9$ $\mathbf{56.0}$ $42.9$
Table 4: Impact of Temporal Module Capacity. Performance on Gen1 dataset with varying ConvLSTM hidden dimensions (M). The rightmost column shows the number of trainable parameters (in millions) in the temporal adaptation module. Minimal performance degradation at M=64 demonstrates efficient adaptation with reduced parameters, similar to LoRA-style approaches.
4.3.3 Parameter Efficiency in Adaptation
Inspired by parameter-efficient techniques like LoRA [17], we investigate the minimal temporal module capacity needed for effective adaptation. Table 4 shows EvRT-DETR-B performance with varying hidden dimensions in the ConvLSTM modules.
Remarkably, reducing hidden dimensions to 64 features (a 4x reduction from our baseline) results in only modest performance degradation ($-0.6\%$ mAP), while requiring just 2.3 M additional parameters—merely $5.4\%$ of the base RT-DETR-B model’s 42.8M parameters. This configuration closely resembles LoRA-style adaptation, achieving substantial performance gains with minimal parameter overhead. Even our most parameter-efficient adaptation (M=64) outperforms all previous state-of-the-art approaches, demonstrating that effective temporal adaptation does not require extensive architectural modifications. Further increasing the hidden dimension to 512 features (42.9 M parameters, nearly doubling the model size) yields only a marginal $0.2\%$ mAP improvement indicating diminishing returns.
5 Conclusions
This work introduces I2EvDet, a framework that efficiently transforms image-based detectors into video-capable models through minimal modifications. Applied to event-based cameras, our approach demonstrates that mainstream object detection architectures can be effectively adapted to temporal domains without specialized redesign. RT-DETR trained on simple event representations achieves performance comparable to specialized methods, while our I2EvDet adapted model EvRT-DETR achieves state-of-the-art results on Gen1 ($+2.3$ mAP) and 1Mpx ($+1.4$ mAP) benchmarks. Our ablation studies provide insights on optimal temporal module design and parameter efficiency. These results suggest that bridging mainstream computer vision and specialized domains can be achieved through targeted adaptation rather than complete architectural redesign, potentially accelerating progress across diverse temporal visual domains.
Acknowledgments
This manuscript has been authored by Brookhaven National Laboratory operated by Brookhaven Science Associates, LLC, under Contract Number DE-SC0012704 with the U.S. Department of Energy. The work was funded by the U.S. Department of Energy, National Nuclear Security Administration, Office of Defense Nuclear Nonproliferation Research and Development.
Appendix A Training Setup
This section details the implementation of our I2EvDet framework as applied to event-based object detection. The training methodology reflects our two-stage adaptation approach: first establishing a strong spatial detector foundation, then incorporating minimal temporal processing while preserving the pre-trained model’s representation capabilities. This implementation demonstrates how mainstream detectors can be efficiently adapted to temporal data with minimal architectural changes
A.1 RT-DETR
Our RT-DETR model is based on the reference PyTorch implementation111commit 5b628eaa0a2fc25bdafec7e6148d5296b144af85 [35, 36]. We explore two RT-DETR configurations: RT-DETR-T (corresponds to RT-DETR-ResNet18) and RT-DETR-B (corresponds to RT-DETR-ResNet50). Our RT-DETR configurations match the reference implementations, except two modifications: (1) we change the number of input channels of the backbones from 3 to 20, and (2) we do not use ImageNet pre-trained backbones. Table 5 summarizes the differences between the RT-DETR-ResNet18 and RT-DETR-ResNet50 configurations.
Configuration RT-DETR-T RT-DETR-B
Backbone ResNet-18 ResNet-50
FPN Features $[128,256,512]$ $[512,1024,2048]$
Encoder Expansion ×$\times 0.5$$\times 1.0$
Decoder Layers 3 6
Table 5: Comparison of the RT-DETR-T and RT-DETR-B Configurations
We train all RT-DETR models for $400{,}000$ iterations using the Adam optimizer [18] with a batch size of 32 and a learning rate of $2\times 10^{-4}$. Similar to the reference RT-DETR training, we maintain an exponential moving average (EMA) of the model weights with a momentum of $0.9999$. Unlike the reference RT-DETR training, we do not use any learning rate schedules, nor do we reduce the learning rate of the backbone relative to the encoder-decoder parts.
Environment.
The RT-DETR training is conducted with PyTorch-2.2.2 and torchvision-0.17.2 on a single NVIDIA RTX A6000 GPU.
A.2 EvRT-DETR
The EvRT-DETR-T and EvRT-DETR-B models extend RT-DETR-T and RT-DETR-B respectively by adding ConvLSTM modules [31]. Our ConvLSTM blocks use a hidden dimension of 256 (matching the encoder feature maps) and a kernel size of 3. The temporal module outputs are integrated into the base model through residual connections with a scaling factor of 1.0, as we found learnable scaling parameters (similar to ReZero [1]) provided no measurable performance benefit on the Gen1 dataset.
The ConvLSTM modules are trained jointly for $200{,}000$ iterations with Adam optimizer and a batch size of 8, while keeping the baseline RT-DETR models frozen. Each batch contains 8 short clips of consecutive frames: 21 frames for Gen1 dataset and 10 frames for 1Mpx dataset. The number of frames per clip follows RVT’s approach [11], except we increase the number of frames for the 1Mpx dataset from 5 to 10 for better performance. Each batch contains 4 randomly sampled clips and 4 consecutive clips from videos.
Unlike the baseline RT-DETR training with a constant learning rate and EMA averaging, we do not apply EMA averaging to the ConvLSTM modules and use a learning rate scheduler instead. For simplicity, we rely on the RVT-inspired one-cycle LR scheduler [32]. Specifically, we use PyTorch’s OneCycleLR implementation of the scheduler, with the following parameters: maximum LR $2\times 10^{-4}$, initial div_factor 20, final div_factor 500, pct_start 0.0005, annealing strategy ”linear”.
The ConvLSTM modules are trained in the same environment as the baseline RT-DETR networks, using a single NVIDIA RTX A6000 GPU.
A.3 Data Augmentation Strategy
While data augmentation is standard practice in computer vision [4, 5], its application to event-based data has been limited [23]. Our adaptation framework treats EBC data as image-like frames, allowing us to leverage established augmentation techniques from mainstream computer vision.
Our augmentation strategy employs a standard chain of geometric transformations and random erasing as summarized in Table 6. This approach bridges conventional image augmentation practices with the unique characteristics of event data, supporting our broader goal of adapting mainstream techniques to event-based vision.
Augmentation Magnitude Probability
Random Horizontal Flip - 0.5
Random Rotation ±$\pm 30^{\circ}$ 0.6
Random Translation ±$\pm 0.5$ 0.6
Random Scale $(0.5,1.5)$ 0.6
Random Shear ±$\pm 30^{\circ}$ 0.6
Random Erasure - 0.4
Table 6: Standard augmentation chain adapted for event-based object detection. Each transformation is applied sequentially with the indicated probability to bridge mainstream vision techniques with event data processing.
Table 6 shows our augmentation strategy, which leverages standard transformations from the torchvision package (0.17.2[21]. Each augmentation is applied independently with its corresponding probability, and when applied, its magnitude is randomly sampled from the specified range. For Random Erasure, we preserve the object labels while erasing image regions, maintaining detection supervision even in partially occluded scenarios.
Temporal Consistency in Augmentations.
Our two-stage adaptation approach requires different augmentation strategies for each stage. For the base detector (RT-DETR) training on individual frames, we apply augmentations independently to each frame following standard practice.
For the temporal adaptation stage, we carefully preserve temporal consistency by applying identical geometric transformations (flip, rotation, scale, translation, shear) across all frames in a video clip, while still allowing per-frame variations through random erasing. This ensures that the temporal module learns meaningful motion patterns rather than artificially induced movements from inconsistent augmentations.
Appendix B Analysis of Base Detector Performance
This section explores key factors that influence the effectiveness of mainstream object detectors on event data, including data augmentation strategies and comparisons across different detector architectures.
B.1 Impact of Augmentation on Base Detector Performance.
When applying mainstream object detectors to event data, appropriate data augmentation becomes crucial for effective domain transfer. Our experiments show that without a proper augmentation strategy, the base detector performance is significantly compromised.
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
RT-DETR-B (ours) $\mathbf{47.6}$ $\mathbf{76.3}$ 49.5
RT-DETR-B (no augs) 38.6 62.8 39.8
RT-DETR-B (-Rotation) 46.8 74.4 48.8
RT-DETR-B (-Scale) 45.6 73.8 47.2
RT-DETR-B (-Translation) 45.0 72.6 46.2
RT-DETR-B (-Shear) $\mathbf{47.6}$ 75.9 $\mathbf{49.7}$
RT-DETR-B (-Erase) 46.8 75.1 48.7
Table 7: Impact of data augmentation techniques on base detector performance for the Gen1 dataset.
Table 7 quantifies the contribution of different augmentation techniques to our adaptation framework. Without augmentations, RT-DETR-B performance drops by 9 mAP points, highlighting their critical role in successful adaptation. Among individual transformations, spatial manipulations (translation and rescaling) provide the largest gains, suggesting that scale and position invariance are particularly important when adapting image detectors to event data. Random rotations and erasure techniques offer moderate improvements, while shear transformations show minimal impact.
B.2 Base Architecture Comparison for Event Data.
To investigate whether our findings about mainstream object detectors on event data extend beyond transformer-based models, we evaluated another popular architecture family - YOLOX [9], a CNN-based detector that has been used in other EBC applications [40, 11].
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
RT-DETR-B (ours) $\mathbf{47.6}$ $\mathbf{76.3}$ $\mathbf{49.5}$
YOLOX-T 36.0 59.9 36.3
YOLOX-S 37.0 61.1 37.6
YOLOX-L 43.1 69.6 44.7
YOLOX-X 43.4 69.7 44.7
Table 8: Performance comparison of different base detector architectures on individual event frames from the Gen1 dataset.
We trained YOLOX models on individual event frames using identical settings to our RT-DETR experiments. As shown in Table 8, YOLOX models achieve respectable performance on the Gen1 dataset, with YOLOX-X reaching 43.4 mAP, which remains below RT-DETR-B’s 47.6 mAP.
These results demonstrate that transformer-based architectures like RT-DETR provide significant advantages for this task compared to CNN-based approaches like YOLOX. The performance differential suggests that our I2EvDet framework benefits from starting with a transformer-based foundation, though the principle of adapting mainstream detectors to event data appears broadly applicable. Testing with the most recent YOLO iterations could be an interesting direction for future work, but was beyond the scope of our current study.
Appendix C Adaptation Considerations for Higher Resolution Event Data (1Mpx)
This section examines specific adaptation factors for the high-resolution 1Mpx dataset, focusing on resolution reduction techniques and temporal context length.
C.1 Resolution Adaptation Strategy
When adapting mainstream detectors to high-resolution event data, appropriate downsampling techniques become critical. For consistency with prior work, we reduce 1Mpx frames from $(720,1280)$ to $(360,640)$, but our investigation reveals that the interpolation method significantly impacts adaptation performance.
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
RT-DETR-B (nearest) 42.3 71.8 42.3
RT-DETR-B (bilinear) $\mathbf{45.2}$ $\mathbf{75.1}$ $\mathbf{46.0}$
RT-DETR-B (bicubic) 43.1 72.2 43.3
Table 9: Impact of interpolation methods on base detector performance for the 1Mpx dataset downsampling.
As shown in Table 9, nearest-neighbor interpolation substantially degrades RT-DETR performance, while bilinear interpolation yields optimal results. Interestingly, bicubic interpolation, despite its theoretical advantages for natural images, proves less effective for event data. These findings align with observations from AEC [23] and highlight the importance of selecting appropriate domain transfer techniques when adapting mainstream vision models to event data.
C.2 Temporal Context Length for Effective Adaptation
The temporal dimension represents a critical aspect of our adaptation framework. While prior work like RVT [11] used 5-frame clips for temporal training, our experiments indicated that expanded temporal context benefits the adaptation process.
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
EvRT-DETR (5 frames) 49.8 80.7 51.8
EvRT-DETR (10 frames) $\mathbf{50.1}$ $\mathbf{80.9}$ $\mathbf{52.1}$
Table 10: Effect of temporal clip length on adaptation performance for the 1Mpx dataset.
Table 10 demonstrates the improvement achieved by extending the temporal window to 10 frames during the adaptation phase. This finding suggests that providing longer temporal context during training allows the model to better capture persistent object representations across the temporal dimension, particularly important for event data where objects may be incompletely represented in shorter time windows.